OpenStudy (anonymous):

25^(2c)=√(5)^(4c+16) solve for c

OpenStudy (anonymous):

c=4

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

25^(2c)=√(5)^(4c+16) 25^(2c)=(5)^(4c+16)]^1/2 squaring both sides we get 25^(4c)=(5)^(4c+16) squaring both sides again we get 25^(8c)=(5)^2(4c+16) 25^(8c)=(25)^(4c+16) here we get (8c)=(4c+16) 4c=16 c=16/4 c=4 ...............ans